After it had cooled the dry marble and test tube weighed 17.56 g. I know it is difficult to see the tick marks for the volume of gas in the video so by the end I collected 35.3 mL of CO2. Calculate the moles of carbon dioxide collected (assume STP and use 4 decimal places)



1
Expert's answer
2021-02-22T05:47:30-0500

CaCO3 + 2HCl = H2O + CO2 + CaCl2

n (CO2) = n (CaCO3)

n (CO2) = n (CaCO3) = V/22.4 = 0.0353/22.4 = 0.002 mol